Assuming that Hund’s rule is violated, the bond order and magnetic nature of the diatomic molecule B2 is
a. 1 and diamagnetic
b. 0 and diamagnetic
c. 1 and paramagnetic
d. 0 and diamagnetic

Grade:upto college level

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hi
B2 = s(1s²) s*(1s²) s(2s²) s*(2s²) p(2px²)
BO=6-4/2=1
And we can see its diamagnetic.
Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty